How many liters of water must be evaporated from 20 L of a 30% salt solution to produce a 50% solution?
Answer by Paul(988)

20(30)=50(20-x)
600=1000-50x
-50x=-400
x=8

Hence, about 8L of water should be evaporated.
PAUL.